OpenStudy (anonymous):

1a. Let x be an arbitrary angle. Verify that (sin x + cos x)^2 = 1 + 2sinxcosx You may assume the Pythagorean identities for arbitrary angles. b. Find an acute angle y such that sin y + cos y = squareroot (2). c. Let x be an arbitrary acute angle. Prove that sin x + cos x >= 1.

OpenStudy (anonymous):

shankvee need ur help!!!!!!!!!

OpenStudy (anonymous):

a) (sin x+ cos x)^2= (sin x)^2 + (cos x)^2 + 2 sin x cos x since (sin x)^2 + (cos x)^2 =1 so the answer is 1 + 2 sin x cos x b) sin y+ cos y =sqrt(2) square 1+2sin y cos y= 2 2sin y cos y=1 2sin y cos y =sin 2y sin 2y=1 2y= 90(degrees) y=45; c)AM > GM (sin x +cos x)/2 >= sqrt(sin x cos x) sin x+ cos >sqrt(2sin 2x) since greatest value of sin 2x =1 so sin x + cos x>sqrt(2) so sin x +cos x>1

OpenStudy (anonymous):

shankvee pls help me!!!!!!

OpenStudy (anonymous):

tejeshwar help you in what?

OpenStudy (anonymous):

Hello shankvee...I dont understand anything you type..

OpenStudy (anonymous):

please can you explain to me some more...

OpenStudy (anonymous):

are you there shankvee

OpenStudy (anonymous):

ok be specific what did you not understand... see (sin x)^2 + (cos x)^2 =1 since; sin x=perpendicular/hypotenuse cos x=base / hypotenuse now square and add since (perpendicular)^2+(base)^2=(hypotenuse)^2

OpenStudy (anonymous):

part b....you said square (1 + 2sin y cos y = 2) Did you mean put it in Square root?

OpenStudy (anonymous):

and part c you said AM > GM..what does that mean?

OpenStudy (anonymous):

are you there sir/maam?

OpenStudy (anonymous):

not sir i am also a student Ok i'll make it simple for you just square the equation in part b and use the identity of part a then you get sin x cos x=1/2 so x=45

OpenStudy (anonymous):

for the third i'll give you an easier answer see (cos x +sin x)^2 = 1+2sin x cos x that means (sin x +cos x)^2 is some value greater than 1 so sin x + cos x is also greater than 1

OpenStudy (anonymous):

THANK YOU!

OpenStudy (anonymous):

I have another problem! Can you help me with it?